OpenStudy (anonymous):

I need an answer checked, I'll fan and medal! // x - 6 / 4 (Fraction) x= -26 The answer I got is 8. Is that correct? If not please explain!

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

i think its right

OpenStudy (anonymous):

but um not shure

OpenStudy (anonymous):

Ok thanks for the help

OpenStudy (anonymous):

wait for owl

OpenStudy (owlcoffee):

All you do is plot the number given and apply the corresponding operations: \[\frac{ x-6 }{ 4 }, x=-26 \iff \frac{ (-26)-6 }{ 4 }=\frac{ -26-6 }{ 4 }\] \[\frac{ -32 }{ 4 }=-8\] So you are correct.

OpenStudy (anonymous):

Wait it's -8?

OpenStudy (owlcoffee):

well, in the number, the sign was the incorrect part.

OpenStudy (owlcoffee):

Yes, it's "-8"

OpenStudy (anonymous):

oh ok thank you

OpenStudy (anonymous):

give owlcoffee the medal cuz he got it right